OpenStudy (anonymous):

If A=x^2 and dx/dt=3, find dA/dt|x= 10 how do i do this because it also says to find dx/dt|x=10 and dA/dt=3

OpenStudy (anonymous):

lol no

OpenStudy (anonymous):

\[A=x^2, A'=2xx'\]

OpenStudy (anonymous):

you are told \(x'=3\) and \(x=10\) so \[A'=2\times 10\times 3=60\]

OpenStudy (anonymous):

since the dx/dt notation sucks,i replaced it by \(x'\) which is easier to write

OpenStudy (anonymous):

hope the answer was clear

OpenStudy (triciaal):

If A=x^2 and dx/dt=3, find dA/dt|x= 10 dA/dx = 2 x dA/dx*dx/dt = dA/dt substitute 2 x * 3 = 10 x = 10/6 = 5/3

OpenStudy (triciaal):

what is the original question?

OpenStudy (anonymous):

A=x^2 is the equation then it says to do the two things

OpenStudy (triciaal):

are you to find dA/dt when x = 10 given that dx/dt = 3 and dA/dt when x = 3

OpenStudy (triciaal):

are those the 2 things?

OpenStudy (anonymous):

yes two separate

OpenStudy (triciaal):

yes that makes sense now

OpenStudy (triciaal):

dA/dx*dx/dt = dA/dt x = 10 and dx/dt = 3 dA/dx = 2 x = 2 * 10 = 20 substituting dA/dt = 20*3 = 60

OpenStudy (anonymous):

but i dont get what the diference si between dx/dt and dA/dt

OpenStudy (triciaal):

dA/dt when x = 3

OpenStudy (triciaal):

dx/dt function x changing with t da/dt change in area with respect to t

OpenStudy (triciaal):

dx/dt = 3 and x = 3 da/dt = da/dx*dx/dt = 2 * 3 * 3 = 18